please help me for these question problem . a test has 20 questions .if lina gets 80percent correct ,how many questions did lina miss to answer correctly
Answer by jim_thompson5910(35256)

There are 20 questions total.
80% of 20 = 80/100*20 = 0.80*20 = 16
So Lina got 16 questions correct

She missed 20-16 = 4 questions

Question: How many questions did she miss?
Answer: 4